OpenStudy (anonymous):

Given: A: {18,6,-3,-12} Determine all elements of set A that are in the solution of the inequality 2/3x +3 <-2x - 7

OpenStudy (anonymous):

can you solve the inequality?

OpenStudy (anonymous):

idk how to im blanking out right now hah

OpenStudy (anonymous):

could you solve this inequality: 3x <2x +1?

OpenStudy (anonymous):

subtract 2x fom both sides

OpenStudy (anonymous):

x<1

OpenStudy (anonymous):

yep, so do same thing here 2/3x +3 <-2x - 7 make that all x's would be in left hand side, how you can do that?

OpenStudy (anonymous):

add -2x so 4/3 x?

OpenStudy (anonymous):

2/3 is fractional ~ 1.333 so adding 2 would be ~3.333 or better \[\frac{2}{3}+2=\frac{2}{3}+\frac{6}{3}=\frac{8}{3}\]

OpenStudy (anonymous):

so what you have left now when you added 2x?

OpenStudy (anonymous):

8/3x+3<-7 then add 3 and you get 8/3<-4?

OpenStudy (anonymous):

if you add 3 to both sides you get 8/3x+3+3<-7+3 8/3x+6<-4 is it correct?

OpenStudy (anonymous):

so it would be 8/3x<-10?

OpenStudy (anonymous):

yes and one more step left, what it is?

OpenStudy (anonymous):

plug in the numbers for x?

OpenStudy (anonymous):

no, if you had 2x<6, how would you solve it?

OpenStudy (anonymous):

divide by 2 on both sides

OpenStudy (anonymous):

so you have to divide there as well by that fraction

OpenStudy (anonymous):

as a decimal its -0.26666666666

OpenStudy (anonymous):

you dont know how to divide fractions? you have to invert second one to change it to the multiplication \[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}\times\frac{d}{c}=\frac{a\times d}{b\times c}\]

OpenStudy (anonymous):

oh yeah sorry im blanking out so much right now

OpenStudy (anonymous):

i got x<-30/8

OpenStudy (anonymous):

yeah, now just convert to decimal for easier comparison and select numbers from A

OpenStudy (anonymous):

okay thank you so much!!!!!!!!!!!!!!!!! you helped so much thank you

OpenStudy (anonymous):

your welcome :)